Equivalent Units of Production and Related Costs The charges to Work in Process—Assembly Department for a...

  1. Equivalent Units of Production and Related Costs

    The charges to Work in Process—Assembly Department for a period, together with information concerning production, are as follows. All direct materials are placed in process at the beginning of production.

    Work in Process-Assembly Department
    Bal., 3,000 units, 60% completed 9,960 To Finished Goods, 69,000 units ?
    Direct materials, 71,000 units @ $1.40 99,400
    Direct labor 173,700
    Factory overhead 67,530
    Bal. ? units, 75% completed ?

    Determine the following:

    a. The number of units in work in process inventory at the end of the period.
    units

    b. Equivalent units of production for direct materials and conversion. If an amount is zero or a blank, enter in "0".

    Work in Process-Assembly Department
    Equivalent Units of Production for Direct Materials and Conversion Costs
    Whole Units Equivalent Units Direct Materials Equivalent Units Conversion
    Inventory in process, beginning
    Started and completed
    Transferred to finished goods
    Inventory in process, ending
    Total units

    c. Costs per equivalent unit for direct materials and conversion. If required, round your answers to the nearest cent.

    Costs Per Equivalent Unit
    Direct Materials $
    Conversion $

    d. Cost of the units started and completed during the period.
    $
Answer #1
a
Number of units in work in process inventory 5000 units =3000+71000-69000
b
Work in Process-Assembly Department
Equivalent Units of Production for Direct Materials and Conversion Costs
Whole Units Equivalent Units Direct Materials Equivalent Units Conversion
Inventory in process, beginning 3000 0 1200 =3000*(1-60%)
Started and completed 66000 66000 66000 =69000-3000
Transferred to finished goods 69000 66000 67200
Inventory in process, ending 5000 5000 3750 =5000*75%
Total units 74000 71000 70950
c
Costs Per Equivalent Unit
Direct Materials 1.40 =99400/71000
Conversion 3.40 =(173700+67530)/70950
d
Units started and completed 66000
X Cost Per Equivalent Unit 4.80 =1.40+3.40
Cost of the units started and completed 316800